# Speech Emotion Recognition with ASR Transcripts: A Comprehensive Study on Word Error Rate and Fusion Techniques

### v1

@@ Abstract
Text data is commonly utilized as a primary input to enhance Speech Emotion Recognition (SER) performance and reliability. However, the reliance on human-transcribed text in most studies impedes the development of practical SER systems, creating a gap between in-lab research and real-world scenarios where Automatic Speech Recognition (ASR) serves as the text source. Hence, this study benchmarks SER performance using ASR transcripts with varying Word Error Rates (WERs) on well-known corpora: IEMOCAP, CMU-MOSI, and MSP-Podcast. Our evaluation includes text-only and bimodal SER with diverse fusion techniques, aiming for a comprehensive analysis that uncovers novel findings and challenges faced by current SER research. Additionally, we propose a unified ASR error-robust framework integrating ASR error correction and modality-gated fusion, achieving lower WER and higher SER results compared to the best-performing ASR transcript. This research is expected to provide insights into SER with ASR assistance, especially for real-world applications.
@@ 1 Introduction
Emotion in speech is conveyed through both what is said (i.e., text) and how it is said (i.e., audio). Recent studies on Speech Emotion Recognition (SER) consistently highlight the benefits of combining textual and acoustic features [1, 2, 3, 4]. Despite the advancements in SER over the years, its practical application in everyday contexts remains limited. One significant challenge is the reliance on human-annotated text, known as gold-standard manual transcripts, in most bimodal SER research. In contrast, transcripts generated by Automatic Speech Recognition (ASR) systems, even state-of-the-art ones, often suffer from high Word Error Rates (WERs), particularly in emotional speech [5]. To this end, the findings from lab settings may not effectively translate to real-world scenarios, impeding the true progress of SER.
To address this gap, it is crucial to investigate the impact of imperfect text data generated by ASR on SER. While previous studies have explored the question of “how ASR performance affects SER”, the findings have been inconsistent and not comprehensive. For instance, Schuller et al. observed that a WER of over 30% resulted in an SER accuracy drop of less than 3% on the FAU Aibo Emotion Corpus [6], while Li et al. reported a nearly 10% accuracy drop with a 30% WER on IEMOCAP [5]. Besides, while multimodal fusion techniques have been proposed and studied extensively to enhance emotion recognition performance, their compatibility with ASR transcripts remains underexplored [7, 8, 9].
In this study, we begin with a benchmark investigation into SER performance utilizing ASR transcripts. First, we generate ASR transcripts from various ASR systems to achieve different WERs. Subsequently, we conduct SER exclusively on these transcripts to assess the impact of ASR errors on text-only SER performance. Finally, we merge audio features and conduct bimodal SER to evaluate whether fusion techniques, effective with ground-truth text, also exhibit efficacy with ASR transcripts (i.e., whether they are ASR error-robust). Following the benchmark study, we propose an ASR error-robust SER framework by integrating ASR error correction with modality-gated fusion to tackle both WER and fusion issues in SER. The contributions of this work are summarized as follows:
$\bullet$ As the first work to benchmark SER performance with ASR transcripts, we offer comprehensive results and novel analysis across diverse corpora, varying WERs, and multiple fusion techniques. This paves the way for SER research with ASR assistance, bringing SER closer to practical applications.
$\bullet$ We propose an ASR error-robust framework that integrates two-stage ASR error correction and dynamic modality-gated fusion, aiming to reduce word errors and mitigate the negative effects of high WER.
2 Related Work: Literature Review
@@
While these studies have highlighted the effectiveness of integrating text features from ASR transcripts, there is still a lack of understanding regarding how WER and fusion techniques impact SER. Therefore, we undertake a benchmark study utilizing diverse ASR models and fusion techniques, conducting SER on various emotion corpora to obtain comprehensive results. Furthermore, drawing from the existing literature, we develop an ASR error-robust framework aimed at mitigating the side effect of WER (Code available [^1] https://github.com/yc-li20/SER-on-WER-and-Fusion).
@@ 3 ASR Models and Emotion Corpora
Emotion corpora. To ensure generalizability, we utilize three corpora: IEMOCAP [19], CMU-MOSI [20], and MSP-Podcast [21], to include diverse speech conditions and various evaluation metrics.
IEMOCAP consists of five sessions of scripted and improvised dialogues conducted in a research lab. Our study focuses on four emotion classes: angry, happy (including excited), neutral, and sad. We exclude utterances with blank transcripts from this corpus, resulting in a total of 5,530 utterances. CMU-MOSI comprises 2,199 monologue video clips sourced from YouTube, each annotated with sentiment scores ranging from -3 to 3. MSP-Podcast contains English speech extracted from podcast recordings, annotated with valence, arousal, dominance scores in the range of 1 to 7, as well as categorical emotion labels. We use its Release 1.11 version and evaluate its performance on the Test1 set. We exclude utterances without an emotion label, resulting in a total of 104,663 utterances.
Following the literature, we calculate Acc4 (four-class accuracy) on IEMOCAP, Acc2 (binary accuracy), Concordance Correlation Coefficient (CCC) on MSP-Podcast, and Acc7 (seven-class accuracy), and Mean Absolute Error (MAE) on CMU-MOSI. Since these metrics are commonly used for respective corpus, we omit detailed explanations.
We initially considered including MELD [22] (TV sitcom data) in our investigation. However, its WERs are almost twice as high as those of the other three corpora (ranging from 30% to 65%). Considering that performing SER using transcripts with such low ASR performance is not feasible in real-world scenarios, we opted to focus only on the other three corpora for our subsequent study. Nonetheless, we provide the WERs of MELD for reference (see Fig. 1).
@@ 4.1 SER Modeling
We employ RoBERTa-base as the text encoder. Given that transcripts generated by models other than Whisper lack punctuation, we remove punctuation from the transcripts of Whisper models to ensure a fair comparison. All letters are lowercased for consistency. A backbone SER model is built for all corpora. Since our goal is not to achieve state-of-the-art performance, the model simply comprises two dense layers: the first encodes RoBERTa output of dimension 768 into hidden states of dimension 128, and the second further encodes it into a dimension of 16. We utilize ReLU as the activation function between the dense layers. In the case of IEMOCAP, we apply one output layer with Softmax activation for classification. For MSP-Podcast and CMU-MOSI, which involve regression tasks, no final output activation is applied. We set the learning rate as 5e-4 for IEMOCAP and CMU-MOSI, and 1e-4 for MSP-Podcast, utilizing the AdamW optimizer. The weight decay is set as 1e-5, and the batch size is 64. For training, we employ five-fold cross-validation on IEMOCAP (100 epochs) and follow the official training/validation/testing sets on CMU-MOSI (100 epochs) and MSP-Podcast (30 epochs). The random seeds are kept consistent across all experiments.
@@ 4.2 Benchmarking SER with WER
Firstly, we present the WERs and corresponding SER performance based on transcripts from different ASR models. From Fig. 1, we observe that:
@@
1) SER performance generally decreases as WER increases. On IEMOCAP and MOSI, there is nearly a 10% accuracy decrease with WERs around 40%, regardless of Acc2, Acc4, or Acc7. However, exceptions exist. For example, on IEMOCAP, Whisper-tiny produces higher accuracy than its neighboring models, even though the WERs are similar. Additionally, W2V960-large-self yields worse accuracy compared to its neighboring models, despite having a relatively lower WER compared to others. The same phenomenon can also been found with CMU-MOSI: Acc2 of W2V100 higher than that of WavLM plus, and of Whisper-large lower than that of Whisper-medium. This might be due to certain words being misrecognized as words that have little effect, or even positively contribute to, their ground-truth emotion labels. Note that the values of our Acc4 are higher than that of [5], possibly because different training models, language models (RoBERTa vs. BERT), and word embeddings (full hidden states vs. pooler output).
@@
Fig. 1: Benchmarking SER performance based on ASR transcripts across corpora. $\uparrow$: higher the better. $\downarrow$: lower the better. Red bold: a higher accuracy than the ground truth.
@@
Fig. 2: Benchmarking SER performance (Acc4) with fusion techniques based on ASR transcripts. Maximum diff: the maximum difference between the Acc4 of ground truth and that of ASR transcripts. Red bold: a higher accuracy than the ground truth.
@@
2) SER is robust to relatively low WER, and in some cases, it is even better with ASR errors. From IEMOCAP, we observe that a WER around 12% barely affects SER performance compared to ground-truth transcripts. Furthermore, on CMU-MOSI, ASR errors can actually enhance SER: transcripts from certain Whisper models outperform the ground truth. Previous studies have confirmed that sentiment analysis is robust to ASR errors [23] in that positive or negative sentiment can contain various emotions. On top of this, we discover that in certain cases (e.g., certain corpora with relatively low WER), ASR errors do not decrease SER performance. In a recent study on speech-based dementia detection, it was found that ASR errors can provide important information for dementia classification [24]. This could be because certain types of ASR errors (e.g., repetitions, insertion of disfluencies) can reveal dementia. However, we do not believe the same holds true for SER. We attribute the phenomenon observed in our study to the same reason as in finding 1: certain words being misrecognized as words that positively contribute to the ground-truth emotions.
@@
3) Different metrics have different sensitivities to WER. On CMU-MOSI, Acc7 exhibits a different pattern from Acc2 and MAE. The variations in Acc2 and MAE are consistent and smooth, while Acc7 appears to be random and lacks a discernible pattern, with almost half of the transcripts (all from Whisper) outperforming the ground truth. This discrepancy may arise from the disparity between the regression model used for training (since MOSI is labeled with continuous values) and the classification metric (where the predicted and ground-truth values are grounded for Acc7), which possibly leads to blurred accuracy and WER-insensitive SER.
@@
4) Different emotion dimensions exhibit varying degrees of robustness to ASR errors. From MSP-Podcast, it is evident that valence, arousal, and dominance exhibit distinct patterns. Firstly, the CCC of valence mirrors the pattern observed in Acc4 for IEMOCAP and Acc2 for CMU-MOSI, suggesting that valence shares similarities with categorical emotion in terms of robustness to ASR errors. Given that valence is similarly defined to sentiment (indicating positive or negative), this alignment is anticipated. Secondly, arousal and dominance do not display a clear correlation with WERs. Since we jointly predict valence, arousal, and dominance, there is a possibility of shared mutual information among the three dimensions, resulting in minimal changes in arousal and dominance. However, our observations align with the the common understanding that valence relies more on text, whereas arousal and dominance rely more on audio [25, 26]. Indeed, when we replaced text features with middle-layer audio features from W2V960-base in the SER model, we obtained CCC_v, CCC_a, and CCC_d values of 0.531, 0.635, and 0.558, respectively, further validating our observations.
@@ 4.3 Benchmarking SER with WER and Fusion
Next, we integrate audio features to investigate the robustness of existing fusion techniques in handling ASR errors in real-world scenarios. As audio is not the primary focus of this study, we simply use the middle-layer audio feature from W2V960-base. For fusion techniques, we employ the following approaches:
@@
• Late fusion: text and audio features are learned independently, and the final decision is determined based on the respective outputs.
@@
For fairness, we keep the backbone SER model used in the previous section unchanged, only modifying the input dimension of the first dense layer to match the output dimension of the hidden states from each fusion model. Additionally, we modify Tensor fusion and MISA to receive bimodal inputs, as they were originally designed for trimodal inputs. Due to limited space, we only present results on IEMOCAP and MSP-Podcast for brevity, as these two are sufficient to cover both categorical and dimensional emotions. Regarding the patterns observed on CMU-MOSI (omitted here), integrating audio did not significantly improve performance, as text alone already yields powerful results. Additionally, the audio in CMU-MOSI does not strongly convey emotion, leading to incongruity and ambiguity, which may pose challenges for SER in certain cases. From Fig. 2, it can be seen that:
@@
1) Fusing audio features can mitigate the negative impact of an increase in WER. The decrease of Acc4 based on WER reaches 10% without fusion on IEMOCAP, but only 4% with fusion. Moreover, with most fusion techniques, the SER performance is even better on transcripts with relatively low WER than on ground truth. These findings confirm the benefits of bimodal SER in real applications, which can be achieved through using a pre-trained ASR model as a base for jointly generating transcripts and audio features [15, 30].
@@
2) There is no optimal WER-robust fusion approach. Due to different amount of input information, the performance of some fusion techniques is not directly comparable. For instance, approaches like Tensor fusion and MISA encode more information due to their incorporation of both unimodal and bimodal features or both modality-invariant and -specific features. However, their effectiveness varies across different corpora. For example, they perform well on IEMOCAP (categorical) but suffer on MSP-Podcast (dimensional), indicating significant inconsistency. Thus, it is hard to identify a single best-performing technique that is universally effective in addressing ASR errors. Nevertheless, the maximum difference between the performance of ground truth and that of ASR transcripts is somewhat indicative of the robustness of certain fusion techniques. We observe that although cross-attention fusion may not produce the best performance, it yields a relatively small maximum difference, suggesting that it is less affected by WER. This observation is reasonable since cross-attention captures the relatedness between bimodal inputs, which changes dynamically once the words have changed, ensuring that the relatedness is least affected by ASR errors.
@@
Since our experimental design has an impact on the results to some extent (e.g., removing punctuation, converting letters to lowercase, excluding samples without corresponding text transcription, modified fusion models), it could inevitably lead to variations and exceptions in some SER performances. However, we believe our overall findings and conclusions remain valid, especially some of them are consistent with existing findings [15, 17, 18].

### v2

@@ Abstract
Text data is commonly utilized as a primary input to enhance Speech Emotion Recognition (SER) performance and reliability. However, the reliance on human-transcribed text in most studies impedes the development of practical SER systems, creating a gap between in-lab research and real-world scenarios where Automatic Speech Recognition (ASR) serves as the text source. Hence, this study benchmarks SER performance using ASR transcripts with varying Word Error Rates (WERs) from eleven models on three well-known corpora: IEMOCAP, CMU-MOSI, and MSP-Podcast. Our evaluation includes both text-only and bimodal SER with six fusion techniques, aiming for a comprehensive analysis that uncovers novel findings and challenges faced by current SER research. Additionally, we propose a unified ASR error-robust framework integrating ASR error correction and modality-gated fusion, achieving lower WER and higher SER results compared to the best-performing ASR transcript. These findings provide insights into SER with ASR assistance, especially for real-world applications.
@@ 1 Introduction
Emotion in speech is conveyed through both what is said (i.e., text) and how it is said (i.e., audio). Recent studies on Speech Emotion Recognition (SER) consistently highlight the benefits of combining textual and acoustic features [1, 2, 3, 4]. Despite the advancements in SER over the years, its practical application in everyday contexts remains limited. One significant challenge is the reliance on human-annotated text, known as gold-standard manual transcripts. In contrast, transcripts generated by Automatic Speech Recognition (ASR) systems, even state-of-the-art ones, often suffer from high Word Error Rates (WERs), particularly in emotional speech [5]. As a result, findings from lab settings may not effectively translate to real-world scenarios, impeding the true progress of SER.
To address this gap, it is crucial to investigate the impact of imperfect text data generated by ASR on SER. While previous studies have explored the question of “how ASR performance affects SER”, the findings have been inconsistent and not comprehensive. For example, Schuller et al. observed that a WER of over 30% resulted in an SER accuracy drop of less than 3% on the FAU Aibo Emotion Corpus [6], whereas Li et al. reported a nearly 10% accuracy drop with a 30% WER on IEMOCAP [5]. Furthermore, while multimodal fusion techniques have been proposed and extensively studied to enhance emotion recognition performance, their compatibility with ASR transcripts remains underexplored [7, 8, 9].
In this study, we begin with benchmarking SER performance using ASR transcripts. We generate ASR transcripts from multiple ASR systems to achieve varying WERs. Subsequently, SER is conducted exclusively on these transcripts to evaluate the impact of ASR errors on text-only SER performance. Following this, we integrate audio features and perform bimodal SER to investigate whether fusion techniques, effective with ground-truth text, also demonstrate efficacy with ASR transcripts (i.e., their resilience to ASR errors). Building on the benchmark study, we propose an ASR error-robust SER framework that integrates ASR error correction with modality-gated fusion to address both WER and fusion challenges in SER. The contributions of this work can be summarized as follows:
$\bullet$ As the first work to benchmark SER performance with ASR transcripts, we examine a wide range of models, varying WERs, and multiple fusion techniques. This sets the stage for advancing SER research with ASR integration, enhancing its practical applications.
$\bullet$ We propose an ASR error-robust framework that integrates two-stage ASR error correction and dynamic modality-gated fusion, aiming to reduce word errors and mitigate the negative effects of high WER on SER performance.
2 Related Work
@@
[…]
@@ 3 ASR Models and Emotion Corpora
Emotion corpora. To ensure generalizability, we utilize three corpora: IEMOCAP [19], CMU-MOSI [20], and MSP-Podcast [21], to include diverse speech conditions and various evaluation metrics for both discrete and continuous emotions.
IEMOCAP consists of five sessions of scripted and improvised dialogues conducted in a research lab. Our study focuses on four emotion classes: angry, happy (including excited), neutral, and sad. We exclude utterances with blank transcripts from this corpus, resulting in a total of 5,500 utterances. CMU-MOSI comprises 2,199 monologue video clips sourced from YouTube, each annotated with sentiment scores ranging from -3 to 3. MSP-Podcast contains English speech extracted from podcast recordings, annotated with valence, arousal, dominance scores in the range of 1 to 7, as well as categorical emotion labels. We use its Release 1.11 version and evaluate its performance on the Test1 set. We exclude utterances without an emotion label, resulting in a total of 104,663 utterances.
Following the literature, we compute Acc4 (four-class accuracy) for IEMOCAP, Concordance Correlation Coefficient (CCC) for MSP-Podcast, and Acc2 (binary accuracy), Acc7 (seven-class accuracy), and Mean Absolute Error (MAE) for CMU-MOSI. Detailed explanations are omitted as these metrics are commonly used for their respective corpora [7, 13].
We initially considered including MELD [22] (TV sitcom data) in our investigation. However, its WERs are nearly double those of the other three corpora, ranging from 30% to 65%. Given that conducting SER using transcripts with such poor ASR performance is impractical in real-world scenarios, we decided to focus on the other three corpora for our subsequent study. Nevertheless, we present the WERs of MELD for reference (see Fig. 1).
@@ 4.1 SER Modeling
We employ RoBERTa-base as the text encoder. Given that transcripts generated by models other than Whisper lack punctuation, we remove punctuation from the transcripts of Whisper models to ensure a fair comparison. All letters are lowercased for consistency. A backbone SER model is built for all corpora. Since our goal is not to achieve state-of-the-art performance, the model simply comprises two dense layers: the first encodes RoBERTa output of dimension 768 into hidden states of dimension 128, and the second further encodes it into a dimension of 16. We use ReLU as the activation function between the dense layers. In the case of IEMOCAP, we apply one output layer with Softmax activation for classification. For MSP-Podcast and CMU-MOSI, which involve regression tasks, no final output activation is applied. We set the learning rate as 5e-4 for IEMOCAP and CMU-MOSI, and 1e-4 for MSP-Podcast, using the AdamW optimizer. The weight decay is set as 1e-5, and the batch size is 64. For training, we employ five-fold cross-validation on IEMOCAP (100 epochs) and follow the official training/validation/testing sets on CMU-MOSI (100 epochs) and MSP-Podcast (30 epochs). The random seeds are kept consistent across all experiments.
@@ 4.2 Benchmarking SER with WER
Firstly, we present the WERs and corresponding SER performance based on transcripts from different ASR models. From Table 1, we observe that:
@@
1) SER performance generally decreases as WER increases. On IEMOCAP and MOSI, there is nearly a 10% accuracy decrease with WERs around 40%, regardless of Acc2, Acc4, or Acc7. However, exceptions exist. For example, on IEMOCAP, Whisper-tiny produces higher accuracy than its neighboring models, even though the WERs are similar. Additionally, W2V960-large-self yields worse accuracy compared to its neighboring models, despite having a relatively lower WER compared to others. The same phenomenon can also be found with CMU-MOSI: the Acc2 of W2V100 is higher than that of WavLM plus, and the Acc2 of Whisper-large is lower than that of Whisper-medium. This might be due to certain words being misrecognized as words that have little effect on or even positively contribute to their ground-truth emotion labels. Note that the values of our Acc4 are higher than those in [5], possibly because of differences in training models, language models (RoBERTa vs. BERT), and word embeddings (full hidden states vs. pooler output).
@@
Table 1: Benchmarking SER performance based on ASR transcripts across corpora. $\uparrow$: higher the better. $\downarrow$: lower the better. Red bold: better performance than the ground truth.
@@
Table 2: Benchmarking SER performance based on ASR transcripts with fusion techniques. Maximum diff: maximum difference between the performance of ground truth and that of ASR transcripts. Red bold: better performance than the ground truth.
@@
2) SER is robust to relatively low WER, and in some cases, it is even better with ASR errors. From IEMOCAP, it is observed that a WER of approximately 12% has minimal impact on SER performance compared to ground-truth transcripts. Moreover, in CMU-MOSI, ASR errors can potentially enhance SER: transcripts generated by specific Whisper models outperform those based on ground truth. Prior research has shown that sentiment analysis remains resilient to ASR errors [23], as positive or negative sentiments can encompass varied emotional states. Additionally, we find that in certain cases (e.g., relatively low WER), ASR errors do not diminish SER performance. Recent research on speech-based dementia detection indicates that ASR errors can offer valuable insights for dementia classification [24], possibly due to specific types of errors (e.g., repetitions, disfluencies) that may reveal indicators of dementia. However, we do not believe this applies similarly to SER. The phenomenon observed in our study may align with our previous finding: certain words being misrecognized as words that positively contribute to the ground-truth emotions.
@@
3) Different metrics have different sensitivities to WER. On CMU-MOSI, Acc7 shows a distinct pattern compared to Acc2 and MAE. Acc2 and MAE demonstrate consistent and smooth variations, whereas Acc7 appears random and lacks a discernible pattern. Interestingly, in nearly half of the transcripts (all from Whisper), Acc7 outperforms the ground truth. This discrepancy may stem from the mismatch between the regression model used during training (since MOSI is labeled with continuous values) and the classification metric applied to Acc7, where predicted and ground-truth values are grounded. This mismatch could potentially blur accuracy and render SER insensitive to WER.
@@
4) Different emotion dimensions exhibit varying degrees of robustness to ASR errors. From MSP-Podcast, it is evident that valence, arousal, and dominance exhibit distinct patterns. Firstly, the CCC of valence mirrors the pattern observed in Acc4 for IEMOCAP and Acc2 for CMU-MOSI, suggesting that valence shares similarities with categorical emotion in terms of robustness to ASR errors. Given that valence is conceptually similar to sentiment (indicating positivity or negativity), this alignment is plausible. Secondly, arousal and dominance do not exhibit a clear correlation with WERs. Since we jointly predict valence, arousal, and dominance, there may be shared information among these dimensions, resulting in minimal changes in arousal and dominance. However, our observations are consistent with the conventional understanding that valence is more influenced by textual content, whereas arousal and dominance are more influenced by audio cues [25, 26, 27]. Indeed, when we replaced text features with middle-layer audio features from W2V960-base in the SER model, we obtained CCC_v, CCC_a, and CCC_d values of 0.531, 0.635, and 0.558, respectively, further validating our observations.
@@ 4.3 Benchmarking SER with WER and Fusion
Next, we integrate audio features to investigate the robustness of existing fusion techniques in handling ASR errors in real-world scenarios. Since audio is not the primary focus of this study, we simply use the middle-layer audio features from W2V960-base. For fusion techniques, we employ the following approaches:
@@
• Late fusion: Text and audio features are learned independently by their respective models, and the final decision is determined based on their outputs.
@@
For fairness, we keep the backbone SER model used in the previous section unchanged, modifying only the input dimension of the first dense layer to match the output dimension of the hidden states from each fusion model. Additionally, we modify Tensor Fusion and MISA to receive bimodal inputs, as they were originally designed for trimodal inputs. Due to limited space, we present results only for IEMOCAP and MSP-Podcast for brevity, as these two corpora are sufficient to cover both categorical and dimensional emotions. Regarding the patterns observed in CMU-MOSI (omitted here), integrating audio did not significantly improve performance, as text alone already yields powerful results. Moreover, the audio in CMU-MOSI does not strongly convey emotion, leading to incongruity and ambiguity, which may pose challenges for SER in certain cases. From Table 2, it can be seen that:
@@
1) Fusing audio features largely mitigates the negative impact of increasing WER. The decrease in Acc4 based on WER reaches 10% without fusion on IEMOCAP, but only 4% with fusion. Moreover, with most fusion techniques, the SER performance is even better on transcripts with relatively low WER than on ground truth. These findings confirm the benefits of bimodal SER in real applications, which can be achieved by using a pre-trained ASR model as a base for jointly generating transcripts and audio features [15, 31].
@@
2) There is no optimal WER-robust fusion approach. Due to different amounts of encoded information, the performances of fusion techniques are not directly comparable. For instance, Tensor Fusion and MISA encode more information by incorporating both unimodal and bimodal features, or both modality-invariant and -specific features. However, their effectiveness varies across different corpora. For example, they perform well on IEMOCAP (categorical) but poorly on MSP-Podcast (dimensional), indicating significant inconsistency. Thus, it is difficult to identify a single best-performing technique that is universally effective in addressing ASR errors. Nevertheless, the maximum difference between the performance of ground truth and that of ASR transcripts is still indicative of the robustness of the fusion techniques. We observe that although cross-attention fusion may not produce the best performance, it yields a relatively small maximum difference, suggesting that it is less affected by WER. This observation is reasonable since cross-attention captures the relatedness between bimodal inputs, which changes dynamically once the words in text have changed, ensuring that the relatedness is least affected by ASR errors.
@@
Since our experimental design impacts the results to some extent (e.g., removing punctuation, converting letters to lowercase, excluding samples without corresponding text transcription, modifying existing fusion models), it could inevitably lead to variations and exceptions in some SER performances. However, our overall findings and conclusions remain valid, especially as some of them are consistent with existing findings [15, 17, 18].